Production priority status

Data production

BulkCorr: Run5 Cu+Cu (QM)

2018-03-09: Was done in SL17i - files are NOT on disk but in HPSS (we may have to verify Cataloging records). Should we assume populating Xrootd with picoDSTs? That previous datasets would be deleted?

2018-03-12: PAC answer

Please restore picoDst to Xrootd


If CuCu 22.4 and 62 data from Run5 have not been reproduced, please reproduce them as well.

Status 2018-03-22:
- As noted, the cu+Cu 200 was done, and we faced Cataloguing issues not yet resolved. Question/Precision: the removal of data could be separated into two issues (1) remove anything from Xrootd/central disk that are from an older production (2) remove from HPSS. We understand the second need holding.

- Lowe Energies: since we started the st_fms, we will not start the low energy production as, during running time, we are lacking tape resources (we do not have enough tape drives to run more than a 2-3 different productions ... and need drives for the MuDST to picoDST conversion). The low energy Cu+Cu in fact cannot be run on the farm at full farm usage (the farm will never saturate). Estimate we could use 30% at most (but again, a choice is to be made: st_fms or Cu+Cu).

Status 2018-03-30: We got confirmation from the PAC (03/23) of replacement old dataset in Xrootd (item #1).The Cu+Cu 200 GeV Cataloguing issue is still unresolved (there has been insufficient time to look into this but some that identified workflow issues that prevent Cataloguing - a solution will be applied whenever time allows).

Spin: Run17 pp510 st_fms

2018-03-09: Has not been scheduled (stand was to understand the issues with the pp 510 based on st_W and st_physics first before launching anything else). Can of course produce “as-is”.

2018-03-12: PAC answered to go ahead "as-is"

Status 2018-03-23: the st_fms was started on 2018-03-17 from a new library release P18ic. As per the progress, the production dataset status page 16.6% were done on the 22nd. ETA: +25 days for the whole sample. Note: we cannot speed up the st_fms as the farm saturation would not be achieved. Currently at 40% of the slots.

Status 2018-03-30: We are now at 56.8% completion. The production rate was feedback as adequate by one PWGC (Carl). Issue with accidental removal of MuDST was noted by the PWGC on 03/26 (Oleg). Automated restore will be set.

Spin/HF/LFSUPC: Run17 pp510 st_physics (on-going)

Restarted/ongoing as discussed in S&C meetings and a few follow-up in the PWG list and S&C noted (8.5% done)

Status as per 2018-03-22: now 20.8% done.
Status 2018-03-30: We are now at 26.9% completion (tape resources becoming an issue - moved two to production but data rate increase is impacting production, priority to st_fms)

LFSUPC: Run17 pp510 st_rp

2018-03-09: Was done as SL18b – requester were informed privately on 2018-03-07.

Status as per 2018-03-22: Although we have asked repeatedly to confirm the PA were informed privately, the PWGC have not provided the courtesy of an answer insofar (+).

Status as per 2018-03-30: We are still waiting for an answer to this simple inquiry.

LFSUPC: Run14 AuAu200 zero-bias

2018-03-09: Not started- pending.

Status as per 2018-03-22: we have not started this production. Estimates of production 1-1.5 weeks. This series is processed twice - once (full) in SL16d with HFT tracking, second time in SL17d without HFT tracking (and selected trigger ID).

Q: We assume this request is for SL17d, no HFT. Please confirm! Also comment on trigger ID.

Possible plan
: We could replace the p+p 500 GeV st_physics with this production. However, we request confirmation of BOTH issues raised by the LSFUPC requests from the PWGC.

Status 2018-03-30: We received no answers related to the chain options (no HFT?) and triggerID (do we use the same?).

HF: Run17 pp510 st_mtd

2018-03-09: See above (same as st_fms)

PWGC answer 2018-03-09:

we prefer to wait for the h+/h- ratio 
puzzle to be understood with the st_physics and st_W data, as stated on 
the summary page.

Status: Pending calibration issue resolution.

HF: Run17 AuAu54 st_mtd

2018-03-09: All Au+Au productions were done in P17ii since 2017/12/15. Not sure what the issue is here.
PWGC answer 208-03-09:

I should have known that the Run17 AuAu 54 st_mtd 
data have already been produced. Thanks.

Topic considered closed.

picoDST production

JetCorr: Run14 Au+Au200 HT no HFT (QM)
Status 2018-03-22: Running, this series will be finished this week (there has been a slow down of production at Dubna, we switched to our online cluster / opportunistic running works well there). Please note again that we do NOT have sufficient tape drive resources at this point in time (restores of MuDST is slow and in wave).
BulkCorr: Run17 Au+Au 54 GeV (QM)
Note: This is a picoDST reproduction to fix an issue with TOF matching.

Status 2018-03-22: The afterburner was integrated (Frank, Daniel, Gene) and available from libraries from SL18c onward. This production has not started yet (implementation was released on the 16th, we cannot start more production in //).

Note: If chain options are provided like the HF PWG, we could start this as soon as Friday the 23rd. The option for the afterburner is simply "btofMatch" (with that option, a geometry MUST be specified i.e. Y2017a in this case)

Note 2018-03-30: Chain options pointed by the PAC on 03/23 as "use the same chain option for the previous production". However, the next wave of jobs for picoDST production was prepared for the p+Au. We will revisit this after the p+Au.


HF: Run15 p+Au 200 st_physics and st_ssdmb (QM)

PWG specified option (TBC): PicoVtxMode:PicoVtxVpdOrDefault, TpcVpdVzDiffCut:6 and specified samples production_pAu200_2015 / st_physics and st_ssdmb  [BEMC important]

Status 2018-03-30: This production has just started (st_physics) and is at 4% completion.

HF: Run16 d+Au 200 st_physics

PWG specified option (TBC): PicoVtxMode:PicoVtxVpdOrDefault, TpcVpdVzDiffCut:6 and samples dAu200_production_2016 / st_physics [BEMC & FMS important]

HF: Run16 Au+Au 200 st_physics

PWG specified option (TBC): PicoVtxMode:PicoVtxVpdOrDefault, TpcVpdVzDiffCut:3 and samples AuAu200_production_2016
AuAu200_production2_2016 / st_physics [BEMC & FMS important]

 HF: Au+Au 200 GeV Run14, st_mtd

PWG specified options (TBC): mtdMatch, y2014a, PicoVtxMode:PicoVtxVpdOrDefault, TpcVpdVzDiffCut:3 and samples AuAu_200_production_2014 && AuAu_200_production_low_2014 && AuAu_200_production_mid_2014 &&
AuAu_200_production_high_2014 / st_mtd

HF: Au+Au 200 GeV Run14, st_physics

PWG  specified options (TBC): PicoVtxMode:PicoVtxVpdOrDefault, TpcVpdVzDiffCut:3 and samples AuAu_200_production_low_2014 AuAu_200_production_mid_2014 / st_physics
[BEMC important]

Status 2018-03-30: This production is done

HF: p+Au 200 GeV, Run 15

PWG specified options (TBC): PicoVtxMode:PicoVtxVpdOrDefault, TpcVpdVzDiffCut:6 and sample production_pAu200_2015 / st_mtd

HF: p+p 200 GeV, Run 15

PWG specified options (TPC): PicoVtxMode:PicoVtxVpdOrDefault, TpcVpdVzDiffCut:6 and sample production_pp200long_2015 &&
production_pp200long2_2015 && production_pp200long3_2015 &&
production_pp200trans_2015 / st_physics and st_ssdmb
[BEMC important]